OpenStudy (anonymous):

The sum of two numbers is 170 and their difference is 26. Find the two numbers. Write the solution as an ordered pair.

OpenStudy (anonymous):

2 To solve, let's call the numbers random variables first. Let's say number 1 is x and number 2 is y. Translate the given statements to equations: The sum of the two numbers is 170: x+ y = 170 The difference in 26: x-y = 26 If we add the two sides of the equation together we get (x+y) + (x-y) = 170+26 2x = 196 x = 98 We have one number, so take another equation to figure out the other numer: x-y = 26 y = x-26, we know x = 98 y = 98-26 = 72 The two numbers are 98 and 72.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

Welcome :)

OpenStudy (anonymous):

what is it as an ordered pair?

OpenStudy (anonymous):

Let the two numbers be a and b then a+b=170 and a−b=26⟹a=98 and b = 72 So ordered pair (98,72)